# On the Effectiveness of ASR Representations in Real-world Noisy Speech Emotion Recognition

### v3

@@ Abstract
In this paper, we propose an efficient approach to noisy speech emotion recognition (NSER). Although conventional NSER methods effectively handle some types of noise, such as stationary noise, they struggle with more complex noise encountered in real-world environments owing to its complexity and unpredictability. To address this issue, we introduce a novel NSER method that leverages automatic speech recognition (ASR) models as noise-robust feature extractors, filtering out non-vocal information from noisy speech. Specifically, we extract intermediate layer representations from the ASR model to capture emotional speech features for the NSER task. In this study, we systematically investigate the effectiveness of ASR representations by comparing them with conventional noise reduction methods and fine-tuned self-supervised learning (SSL) approaches. We categorize ASR representations into Encoder and Decoder groups and analyze the last layer, the mean of all layers, and a proposed adapter method. Additionally, we assess the robustness of ASR representations under various noise intensities and types, explore their effects across different modalities, and compare the model’s performance with ASR transcripts. Our experimental results demonstrate that (1) our method significantly outperforms conventional noise reduction and SSL approaches, and (2) our method optimizes ASR representations with a proposed adapter-based method, resulting in significant gains for NSER tasks under various noise intensities. However, (3) its performance diminishes with increasing noise intensity, particularly in complex scenarios involving human speech noise, and (4) this decline mirrors the degradation observed in ASR performance as noise intensity increases. Notably, (5) it also surpasses text-based approaches using ASR or ground-truth transcription of noisy speech and (6) exhibits robust performance under cross-lingual conditions compared with mainstream self-supervised representations.
Index Terms:
Speech Emotion Recognition, Automatic Speech Recognition, Self-Supervised Learning.I Introduction
@@
Speech emotion recognition (SER) has attracted considerable attention over the past two decades, particularly within the domain of human-computer interaction [1, 2, 3]. SER has enabled a variety of applications, fundamentally transforming customer service interactions by allowing chatbots to provide empathetic and contextually appropriate responses [4, 5]. Additionally, SER has been integrated into in-vehicle dashboard systems, improving driver experience and safety through real-time understanding and responsiveness to emotional states [6, 7]. Moreover, the utility of this technology extends to speech-to-speech translation platforms, enhancing the accuracy and naturalness of cross-linguistic communication by incorporating emotional context [8, 9]. Despite these promising applications, SER systems face a significant challenge: the prevalence of unknown noise in various real-world scenarios. This noise severely hampers the performance of SER models, preventing them from effectively distinguishing and accurately recognizing emotions amidst environmental disturbances. Such limitations pose a considerable barrier to the widespread adoption and effective implementation of SER across diverse domains [10, 11].
@@
Empirical investigations have demonstrated the efficacy of established methodologies in mitigating common noise sources, such as white Gaussian noise, in speech-related tasks [12, 13, 14, 15, 16, 17]. However, translating such successes to real-world settings faces challenges posed by a distinct category of noise, which includes ambient sounds such as the clicking of high-heeled shoes or the abrupt knocking of doors. Unlike Gaussian noise, these real-world noises exhibit non-Gaussian distributions and possess inherent unpredictability, complicating their effective reduction or elimination through conventional noise reduction techniques [10, 18]. Consequently, the performance of SER systems, already vulnerable to noise, faces greater limitations in accurately recognizing emotional cues amidst such disturbances [19]. Therefore, effectively adapting SER systems to tackle these real-world noise challenges is crucial for enhancing their robustness and applicability across diverse contexts.
@@
Capitalizing on this intrinsic capability, we propose an innovative approach to address the challenge of noisy speech emotion recognition (NSER) in real-world scenarios. In our framework, ASR representations play a pivotal role in capturing fundamental vocal cues that underpin emotional expressions. By leveraging the inherent robustness of ASR-derived features to noise, we aim to mitigate the harmful effects of environmental disturbances on emotion recognition accuracy. In this study, we explore the impact of ASR representations on NSER by comparing them with conventional noise reduction techniques, specifically CONVTAS [24] and DCCRN [25], as well as fine-tuning methods based on self-supervised learning (SSL) models such as Wav2vec 2.0 [26], HuBERT [27], and WavLM [28]. A comprehensive analysis is conducted to assess the effectiveness of various ASR representations. This includes a comparative evaluation of different levels of encoder and decoder representations within the ASR model for NSER. We also employ a layer adapter method to integrate representations from different levels. The robustness of ASR representations is further examined under various noise intensities by testing our model across different signal-to-noise ratio (SNR) conditions. We investigate the correlation between ASR performance and NSER performance by testing the ASR model under various noise intensities and types. To evaluate the effectiveness of ASR representations in NSER, we assess their performance across different modalities. Finally, we test the robustness of our proposed model in cross-lingual NSER scenarios.
@@
Introduction of a novel approach to addressing real-world noise challenges in NSER by utilizing ASR.
Proposal of an innovative method for leveraging ASR encoder and decoder representations, specifically tailored for NSER.
@@
The remaining sections of this paper are organized as follows: In Section II, we provide an overview of NSER and the application of ASR in SER, forming the foundation for the proposed approach. In Section III, we present the proposed method and model structure. In Section IV, we describe the three databases used in this study. In Section V, we outline the experimental setup, including the control group design, implementation details, and evaluation metrics. In Section VI, we present the experimental results, validating the effectiveness of the proposed model. Finally, in Section VII, we provide a discussion and suggest directions for future research.
@@ II-A Noisy Speech Emotion Recognition
In real-world scenarios, various types of noise present themselves, each with distinct characteristics and sources [29]. Gaussian white noise is a typical example, representing a fundamental manifestation of signal randomness. It is characterized as stationary noise with a uniform distribution of power spectral density across all frequencies, symbolizing an essential element of signal variability [30]. Conversely, impulse noise consists of sudden, high-amplitude bursts of pulses or spikes, often arising from equipment malfunctions or electromagnetic interference incidents [31]. Additionally, environmental noise encompasses a broad spectrum of background disturbances from the surrounding environment, including sources such as vehicular noise, industrial machinery operations, and human activities, all contributing to real-world noise [32].
@@
To address the multifaceted challenges posed by the diverse noise patterns encountered in real-world scenarios, numerous research initiatives have been undertaken to mitigate the detrimental impact of noise on the accuracy of speech-related tasks, particularly in the domain of NSER. These endeavors typically employ three primary strategies: signal-level, feature-level, and model-level interventions, as elucidated by Tiwari et al. [12]. For example, Pandharipande et al. [13] proposed an unsupervised method that uses a front-end voice activity detector to selectively extract frames with higher SNRs from spoken utterances, aiming to enhance accuracy under noisy conditions by improving signal quality at the signal level. Similarly, Vasquez-Correa et al. [33] proposed a method for emotion recognition using wavelet packet transform features. They classified negative emotions and distinguished them from neutral and positive states by analyzing voiced and unvoiced segments. Features such as log energy and mel-frequency cepstral coefficients were calculated, and the method was tested on the Berlin Emotional Database and the eNTERFACE05 Database under various noise conditions. They also evaluated two speech enhancement techniques, noting improved accuracy in noisy environments. At the feature level, Chenchah and Lachiri [15] introduced a novel approach involving MFCC-shifted-delta-cepstral coefficients. On the other hand, Sekkate et al. [16] proposed an approach combining baseline mel-frequency cepstral coefficients, discrete wavelet transform-derived MFCCs, and pitch-based features to create a feature set that effectively captures relevant information in a relatively low-dimensional space. At the model level, Tiwari et al. [12] devised multi-conditioning and data augmentation using an utterance-level parametric generative noise model. This strategy aims to encompass the entire noise space in the mel-filterbank energy domain, rendering the model robust against unseen noise conditions. Additionally, Nam and Lee [17] introduced a novel approach employing a cascaded denoising CNN architecture for SER under noisy conditions. This architecture consists of two stages: DnCNN for denoising using residual learning and CNN for subsequent classification at the model level.
@@ II-B Automatic Speech Recognition for Speech Emotion Recognition
ASR is a critical tool for transcribing spoken language into written text, with broad applicability across various real-world domains [34]. Significant progress has been made in the field of SER through the use of ASR-generated transcripts. Yoon et al. [35] pioneered a deep dual recurrent encoder model that simultaneously leveraged both audio signals and text data from the Google Cloud Speech API. Subsequently, Sahu et al. [36] used transcripts from two commercial ASR systems for a bimodal SER approach (audio and text), noting a relative decline in unweighted accuracy compared with ground-truth transcripts. Li et al. [37] introduced a temporal alignment mean-max pooling mechanism to capture subtle, fine-grained emotions within utterances, as well as a cross-modality excitement module to make sample-specific adjustments to embeddings. Santoso et al. [38] proposed a confidence measure to adjust the importance weights of ASR transcripts based on the likelihood of recognition errors in each word, effectively mitigating the impact of ASR errors on SER performance. Wu et al. [39] developed a dual-branch model to enhance robustness against ASR errors, featuring a time-synchronous branch that combines speech and text modalities, alongside a time-asynchronous branch that integrates sentence text embeddings from contextual utterances. Shon et al. [40] generated pseudo labels for ASR transcripts in semi-supervised speech sentiment analysis. On the basis of human perception mechanisms, Li et al. [41] proposed the hierarchical attention fusion of audio features, ASR hidden states, and ASR transcripts, achieving SER performance comparable to that using ground-truth text. Lin and Wang [42] explored complementary semantic information from audio to mitigate ASR errors, employing an attention mechanism to calculate weighted acoustic representations on the basis of ASR hypotheses. He et al. [43] introduced two auxiliary tasks—ASR error detection and correction—to improve the semantic coherence of ASR text, and proposed a novel multimodal fusion method to learn shared representations across modalities. Finally, Feng and Narayanan [44] fused audio with ASR transcripts from an advanced ASR model, demonstrating that ASR-generated outputs deliver competitive SER performance relative to ground-truth transcripts.
ASR models inherently prioritize capturing human vocals while effectively filtering out non-vocal components from speech [23]. Building on the efficacy of ASR models in capturing human speech, in this study, we propose leveraging ASR representations as robust features for emotion recognition.
@@ III-B ASR Embedding Module
Different from most existing models that recognize emotional states using SSL and text features, the proposed model utilizes ASR representations to offer advancements in emotion recognition from several perspectives. Whereas SSL presents the advantage of leveraging extensive unsupervised data for model training, recent ASR systems such as Whisper have primarily benefited from supervised training with substantial labeled datasets. Generally, when ample labeled data is available, supervised learning methods tend to exhibit superior performance. Additionally, ASR representations can capture more speech temporal structures than text features. Therefore, we opt for ASR as our feature extractor.
@@ IV-A Database
In this section, we introduce the databases used in this study: the MELD dataset, which contains mixed real-world scenario recordings; the IEMOCAP dataset, offering clean recordings; and the CASIA dataset, designed for investigating cross-lingual scenarios. The emotional distribution across these datasets is illustrated in Fig. 2.
@@
The MELD dataset is an extension of a sentiment-focused dataset, comprising 1,433 dialogues from the American drama “Old Friends,” totaling 13,708 sentences and incorporating various data types, including video, text, and speech [51]. Various real-world noises are recorded within the MELD dataset, such as car horn sounds, knocking noises, plate rattling sounds, and more. Each dialogue in MELD is annotated with seven emotions: anger, disgust, sadness, happiness, neutrality, surprise, and fear. Additionally, the dataset provides standard training, validation, and test sets to facilitate comparative experimental results. Our study focused on four emotions: happiness, anger, sadness, and neutrality.
@@
For this study, we focus on four categorical emotions: neutral, happiness, sadness, and anger, aligning with experimental protocols used in previous research [49, 50, 53]. Additionally, we combine happiness and excitement into a single category. Given that the IEMOCAP recordings were primarily captured in clean environments, we introduce noise from the DEMAND corpus to simulate real-world noisy conditions. The DEMAND corpus includes six noise categories: Street, Domestic, Office, Public, Nature, and Transportation [54]. For our study, we selected Office and Nature noises to represent distinct environmental conditions.
2.1 IEMOCAP-Environment
To simulate real-world scenarios, we combined the IEMOCAP data with the Nature category from the DEMAND corpus. Specifically, we used the NPARK segment of the Nature data, recorded in a bustling city park, to augment the IEMOCAP dataset.

### v4

# A Comprehensive Study on the Effectiveness of ASR Representations for Noise-Robust Speech Emotion Recognition

@@ Abstract
In this paper, we propose an efficient noise-robust approach to noisy speech emotion recognition (NSER). Although conventional NSER methods effectively handle some types of noise, such as stationary noise, they struggle with more complex noise encountered in realistic acoustic environments owing to its complexity and unpredictability. To address this issue, we introduce a novel NSER method that leverages automatic speech recognition (ASR) models as noise-robust feature extractors, filtering out non-vocal information from noisy speech. Specifically, we extract intermediate layer representations from the ASR model to capture emotional speech features for the NSER task. In this study, we conduct a comprehensive evaluation of ASR representations against traditional high-level statistical function (HSF) features, noise reduction approaches, mainstream NSER models, and fine-tuned self-supervised learning (SSL) methods. We analyze encoder and decoder groups through both single-layer and fusion-based multi-layer representations, further investigating the layer-wise performance and robustness of encoder and decoder modules under different noise types and intensities. Additionally, we explore their correlation with ASR performance, assess cross-modal performance using ASR transcripts, and examine cross-lingual generalization. Our experimental results demonstrate that (1) the proposed method outperforms HSF features, noise reduction approaches, mainstream NSER models, and SSL methods; (2) the adapter-based method further enhances ASR representations; (3) higher encoder and lower decoder layers (e.g., in Whisper) yield the best individual performance, while combining encoder and decoder layers across all depths achieves the strongest results; (4) performance decreases as noise intensity increases, especially with human speech noise; (5) this decline mirrors the degradation observed in ASR performance as noise intensity increases; (6) our approach surpasses transcript-based methods using ASR or ground-truth transcriptions; and (7) it achieves robust cross-lingual performance compared with mainstream SSL representations.
I Introduction
@@
Speech emotion recognition (SER) has attracted considerable attention over the past two decades, particularly within the domain of human-computer interaction [58, 12, 6]. SER has enabled a variety of applications, fundamentally transforming customer service interactions by allowing chatbots to provide empathetic and contextually appropriate responses [13, 19]. Additionally, SER has been integrated into in-vehicle dashboard systems, improving driver experience and safety through real-time understanding and responsiveness to emotional states [23, 78]. Moreover, the utility of this technology extends to speech-to-speech translation platforms, enhancing the accuracy and naturalness of cross-linguistic communication by incorporating emotional context [2, 14]. Despite these promising applications, SER systems face a significant challenge: the prevalence of unknown noise in realistic acoustic environments. This noise severely hampers the performance of SER models, preventing them from effectively distinguishing and accurately recognizing emotions amidst environmental disturbances. Such limitations pose a considerable barrier to the widespread adoption and effective implementation of SER across diverse domains [77, 52].
@@
Empirical investigations have demonstrated the efficacy of established methodologies in mitigating common noise sources, such as white Gaussian noise, in speech-related tasks [69, 46, 44, 10, 59, 45]. However, translating such successes to realistic acoustic environments remains challenging due to a distinct category of noise, including ambient sounds such as the clicking of high-heeled shoes or abrupt door knocking. Unlike Gaussian noise, these environmental noises exhibit non-Gaussian distributions and inherent unpredictability, complicating their effective reduction or elimination through conventional noise reduction techniques [77, 26]. Consequently, the performance of SER systems, already vulnerable to noise, faces greater limitations in accurately recognizing emotional cues under such disturbances [4]. Therefore, effectively adapting SER systems to complex acoustic environments is crucial for enhancing their robustness and applicability across diverse contexts.
@@
Capitalizing on this intrinsic capability, we propose an innovative approach to address the challenge of noisy speech emotion recognition (NSER) in realistic acoustic environments. In our framework, ASR representations play a pivotal role in capturing fundamental vocal cues that underpin emotional expressions. By leveraging the inherent robustness of ASR-derived features to noise, we aim to mitigate the harmful effects of environmental disturbances on emotion recognition accuracy. First, we compare ASR representations with several strong baselines, including traditional high-level statistical function (HSF) features [57, 54, 53, 55, 56, 15], conventional noise reduction methods [41, 28], mainstream NSER models [39, 38], and fine-tuned self-supervised learning (SSL) models [3, 27, 9]. Second, we examine the effectiveness of single-layer and fusion-based multi-layer ASR representations, where the single-layer setting adopts the last layer or the best-performing layer among all layers, and the fusion setting includes simple fusion (mean pooling), parameterized fusion (weighted sum and scalar mixing), and advanced fusion (attention-based pooling). Third, we conduct a layer-wise analysis to compare the performance of encoder and decoder representations, further contrasting them with SSL features. Fourth, we assess robustness to different noise intensities by testing under various signal-to-noise ratio (SNR) conditions. Fifth, we investigate the correlation between ASR and NSER performance by analyzing their degradation trends across noise conditions. Sixth, we evaluate the effectiveness of ASR representations in NSER by assessing their performance across different modalities. Finally, we assess the generalization of ASR representations by testing their effectiveness in cross-lingual NSER scenarios.
@@
Introduction of a novel ASR-based approach for noise-robust NSER in realistic acoustic environments.
Proposal of a method for leveraging ASR encoder and decoder representations for NSER.
Conducting a systematic layer-wise analysis of ASR models, which reveals that higher encoder layers and lower decoder layers (e.g., Whisper) contribute most effectively to NSER.
@@
The remaining sections of this paper are organized as follows: In Section II, we provide an overview of NSER and the application of ASR in SER, forming the foundation for the proposed approach. In Section III, we present the proposed method and model structure. In Section IV, we describe the three databases used in this study. In Section V, we outline the experimental setup, including the control group design, implementation details, and evaluation metrics. In Section VI, we present the experimental results, validating the effectiveness of the proposed model. Finally, in Section VII, we discuss findings and outline future work.
@@ II-A Noisy Speech Emotion Recognition
In realistic acoustic environments, various types of noise arise, each with distinct characteristics and sources [30]. Gaussian white noise is a typical example, representing a fundamental manifestation of signal randomness. It is characterized as stationary noise with a uniform distribution of power spectral density across all frequencies, symbolizing an essential element of signal variability [42]. Conversely, impulse noise consists of sudden, high-amplitude bursts of pulses or spikes, often arising from equipment malfunctions or electromagnetic interference incidents [25]. Additionally, environmental noise encompasses a broad spectrum of background disturbances from the surrounding environment, including sources such as vehicular noise, industrial machinery operations, and human activities, all contributing to environmental noise [43].
@@
To address the multifaceted challenges posed by diverse noise patterns in realistic acoustic environments, numerous studies have been conducted to mitigate the detrimental impact of noise on the accuracy of speech-related tasks, particularly in NSER. These endeavors typically employ three primary strategies: signal-level, feature-level, and model-level interventions, as elucidated by Tiwari et al. [69]. For example, Pandharipande et al. [46] proposed an unsupervised method that uses a front-end voice activity detector to selectively extract frames with higher SNRs from spoken utterances, aiming to enhance accuracy under noisy conditions by improving signal quality at the signal level. Similarly, Vasquez-Correa et al. [70] proposed a method for emotion recognition using wavelet packet transform features. They classified negative emotions and distinguished them from neutral and positive states by analyzing voiced and unvoiced segments. Features such as log energy and mel-frequency cepstral coefficients were calculated, and the method was tested on the Berlin Emotional Database and the eNTERFACE05 Database under various noise conditions. They also evaluated two speech enhancement techniques, noting improved accuracy in noisy environments. At the feature level, Chenchah and Lachiri [10] introduced a novel approach involving MFCC-shifted-delta-cepstral coefficients. On the other hand, Sekkate et al. [59] proposed an approach combining baseline mel-frequency cepstral coefficients, discrete wavelet transform-derived MFCCs, and pitch-based features to create a feature set that effectively captures relevant information in a relatively low-dimensional space. At the model level, Tiwari et al. [69] devised multi-conditioning and data augmentation using an utterance-level parametric generative noise model. This strategy aims to encompass the entire noise space in the mel-filterbank energy domain, rendering the model robust against unseen noise conditions. Additionally, Nam and Lee [45] introduced a novel approach employing a cascaded denoising CNN architecture for SER under noisy conditions. This architecture consists of two stages: DnCNN for denoising using residual learning and CNN for subsequent classification at the model level. More recently, Liu et al. [39] introduced a multi-level knowledge distillation framework that transfers robust knowledge across network depths, thereby enhancing SER resilience under noisy conditions. Building on this, Liu et al. [38] proposed an adaptive emotion denoising diffusion model combined with an iterative confidence learning strategy, achieving improved performance in noisy environments.
@@ II-B Automatic Speech Recognition for Speech Emotion Recognition
ASR is a critical tool for transcribing spoken language into written text, with broad applicability across various diverse application domains [67]. Significant progress has been made in the field of SER through the use of ASR-generated transcripts. Yoon et al. [76] pioneered a deep dual recurrent encoder model that simultaneously leveraged both audio signals and text data from the Google Cloud Speech API. Subsequently, Sahu et al. [50] used transcripts from two commercial ASR systems for a bimodal SER approach (audio and text), noting a relative decline in unweighted accuracy compared with ground-truth transcripts. Li et al. [33] introduced a temporal alignment mean-max pooling mechanism to capture subtle, fine-grained emotions within utterances, as well as a cross-modality excitement module to make sample-specific adjustments to embeddings. Santoso et al. [51] proposed a confidence measure to adjust the importance weights of ASR transcripts based on the likelihood of recognition errors in each word, effectively mitigating the impact of ASR errors on SER performance. Wu et al. [73] developed a dual-branch model to enhance robustness against ASR errors, featuring a time-synchronous branch that combines speech and text modalities, alongside a time-asynchronous branch that integrates sentence text embeddings from contextual utterances. Shon et al. [63] generated pseudo labels for ASR transcripts in semi-supervised speech sentiment analysis. Feng and Narayanan [16] fused audio with ASR transcripts from an advanced ASR model, demonstrating that ASR-generated outputs deliver competitive SER performance relative to ground-truth transcripts.
Beyond transcripts, several studies have investigated ASR model representations directly for SER. Lu et al. [40] employed pre-trained features from end-to-end ASR models for SER, showing that hidden states from ASR encoders integrate both acoustic and linguistic information. Yeh et al. [75] proposed a factorized adaptation approach that learns disentangled speaker and linguistic representations from end-to-end ASR models, achieving effective transfer to emotion recognition tasks. Li et al. [36] proposed the hierarchical attention fusion of audio features, ASR hidden states, and ASR transcripts, achieving SER performance comparable to that using ground-truth text. Lin and Wang [37] explored complementary semantic information from audio to mitigate ASR errors, employing an attention mechanism to calculate weighted acoustic representations on the basis of ASR hypotheses. Gao et al. [18] further designed a multi-level extraction framework that integrates acoustic and semantic information, where semantic representations are derived from ASR outputs and combined with acoustic features through cross-attention and gating. He et al. [24] introduced two auxiliary tasks, ASR error detection and correction, to improve the semantic coherence of ASR text, and proposed a novel multimodal fusion method to learn shared representations across modalities.
However, most prior studies have focused on clean data conditions and overlooked an intrinsic property of ASR models: while encoders capture rich acoustic-linguistic cues, decoders also encode complementary semantic and contextual information through their language modeling role [49, 65]. Such information is particularly valuable for emotion recognition, where lexical and prosodic cues jointly contribute to performance [72]. Meanwhile, the potential utility of decoder-derived representations has received little attention, and the relative robustness of encoder versus decoder features under diverse noise conditions has not been systematically examined. To address this gap, our work investigates both encoder and decoder representations as noise-robust features for NSER, providing the first comprehensive analysis of their effectiveness and complementarity in realistic acoustic environments.
@@ III-B ASR Embedding Module
Different from most existing models that recognize emotional states using SSL and text features, the proposed model utilizes ASR representations to offer advancements in emotion recognition from several perspectives. Whereas SSL presents the advantage of leveraging extensive unsupervised data for model training, recent ASR systems such as Whisper have primarily benefited from supervised training with substantial labeled datasets. Generally, when sufficient labeled data is available, supervised learning methods tend to exhibit superior performance. Additionally, ASR representations can capture more speech temporal structures than text features. Therefore, we opt for ASR as our feature extractor.
@@ IV-A Database
In this study, we employ three emotional speech corpora. The MELD dataset, which contains mixed recordings from realistic acoustic environments, has been widely used for NSER research [74, 31]. The IEMOCAP dataset provides relatively clean and controlled recordings, and the CASIA dataset is used to explore cross-lingual scenarios. The emotional distribution across these datasets is illustrated in Fig. 2.
@@
The MELD dataset is an extension of a sentiment-focused dataset, comprising 1,433 dialogues from the American drama “Old Friends,” totaling 13,708 sentences and incorporating various data types, including video, text, and speech [48]. Various environmental noises are recorded within the MELD dataset, such as car horn sounds, knocking noises, plate rattling sounds, and more. Each dialogue in MELD is annotated with seven emotions: anger, disgust, sadness, happiness, neutrality, surprise, and fear. Additionally, the dataset provides standard training, validation, and test sets to facilitate comparative experimental results. Our study focused on four emotions: happiness, anger, sadness, and neutrality.
@@
For this study, we focus on four categorical emotions: neutral, happiness, sadness, and anger, aligning with experimental protocols used in previous research [35, 17]. Additionally, we combine happiness and excitement into a single category. Given that the IEMOCAP recordings were primarily captured in clean environments, we introduce noise from the DEMAND corpus to simulate realistic acoustic conditions. The DEMAND corpus includes six noise categories: Street, Domestic, Office, Public, Nature, and Transportation [68]. For our experiments, we selected Office (representing indoor environments), Nature, and Transportation (representing outdoor environments) to construct distinct noisy scenarios.